OpenStudy (anonymous):

cos(x)−2cos(x) sin(x)=0

OpenStudy (anonymous):

Find value of x cos(x)−2cos(x) sin(x)=0

zepdrix (zepdrix):

\[\Large\bf\sf \cos x-2 \cos x \sin x=0\]Factor a cos x out of each term,\[\Large\bf\sf \cos x(1-2\sin x)=0\]Apply the Zero Factor Property, setting each factor equal to zero,\[\Large\bf\sf \cos x=0,\qquad\qquad 1-2\sin x=0\]Solve for x in each case.